# Real-Time Moving Flock Detection in Pedestrian Trajectories Using Sequential Deep Learning Models

### v1

@@ 3.2.1 Time-Bin Classification
[…]
@@ 3.3.1 Input and Output
3.3.2 Algorithm Description
The algorithm operates as follows:
@@ 4.1.1 Dataset
The group files contain annotations for the groups present on a given day. Only pedestrians in groups are listed, while those walking alone are excluded. Each row represents a single tracked pedestrian within a group and includes the following fields (space-separated values): $PEDESTRIAN-ID$, $GROUP-SIZE$, $PARTNER-ID-1$… (list of IDs of all other pedestrians in the group), $NUMBER-OF-INTERACTING-PARTNERS$, and $INTERACTING-PARTNER-ID-1$… (list of all socially interacting partners).
For the binary classification model training, we select only data where $GROUP-SIZE$ = 2, focusing on pairs of pedestrians from a single day. The pre-trained model is then tested on data from different days, ensuring that the model is evaluated on entirely unseen data.
@@ 4.1.2 Deep Learning Model Preparation
In accordance with the definition of a flock, we extract features that capture spatial and behavioral interactions between pedestrians. These include the center of the pair, the inter-distance between pedestrians, and absolute differences in motion properties such as velocity, motion angle, and facing angle. Additionally, we incorporate trajectory-based similarity metrics like Dynamic Time Warping (DTW) to enhance the model’s ability to distinguish between temporary co-movement and sustained flocking behavior.
@@
1 centerX X-coordinate of the center between the two pedestrians in the pair.
2 centerY Y-coordinate of the center between the two pedestrians in the pair.
@@ 4.1.3 Multi-agent group data preparation
The total number of bins in the dataset is calculated to be 255. These bins serve as the basis for further analysis, including the detection of flocking behavior and the execution of pairwise interactions, which are crucial for understanding collective pedestrian dynamics.
@@ 4.3.1 Binary Classification Accuracy by Data Length and Models
Figure 3: Sequential models accuracy for different sequences
@@
4.3.2 RunTime Comparison
@@
The runtime comparison graph indicates that the RNN model has the shortest training time, making it the fastest among the three models due to its simpler structure. In contrast, the LSTM model has the longest training time across all sequence lengths, likely due to its more complex memory mechanisms.
The training time also varies with sequence length. Interestingly, the fastest training occurred at sequence lengths 100 and 300, while sequence length 150 took the longest to train. This suggests that certain sequence lengths may introduce computational inefficiencies.
@@
  • RNN is the fastest model across all sequence lengths due to its simpler architecture.
  • LSTM has the longest training time, peaking at 57.1 minutes for sequence length 150.
  • Transformer runtimes closely follow LSTM but are slightly shorter in most cases.
  • Sequence lengths 100 and 300 resulted in the shortest runtimes across all models, while 150 took the longest to train.
These results highlight the trade-offs between model complexity and training efficiency. While LSTM and Transformers are more powerful for sequence learning, they require significantly longer training times than RNNs. Choosing the right sequence length can also impact training efficiency, with some lengths (e.g., 100, 300) being computationally favorable.
@@ 4.3.6 Sample Screens for flock detection
(c) Detected flock size with 7.
Figure 7: Sample flock detection screen steps with 4 different sizes
@@ 5 Discussion
In this section, we discuss the results of our flock detection algorithm and highlight key observations regarding the impact of data sequence length and the dependency of features used in the model.
@@
The current feature set for the binary classification model, consisting of center coordinates (centerx, centery), inter-agent distance, absolute values of the timestamp, velocity, motion angle, and face angle, has shown promising results in detecting pairs. With a total of 8 features, the model achieved high performance. Interestingly, in some instances, reducing the number of features to 6 (excluding centerx and centery) still yielded similar accuracy results. This suggests that the model can effectively detect pairs based on more dynamic features, such as inter-agent distance, velocity, and motion angles. The less significant contribution of static features like center coordinates in certain instances could inform future efforts to optimize the feature set, potentially simplifying the model without sacrificing performance. In practice, reducing the feature space could improve computational efficiency while maintaining high accuracy, especially for real-time applications.

### v2

@@ 3 Methodology
In this context, a flock is formally defined as a subset of pedestrians whose trajectories exhibit sustained spatial proximity and coherent motion patterns over a contiguous temporal interval. The spatial constraint is characterized by bounded inter-pedestrian distances and consistent relative orientations, while the temporal constraint requires persistence of these spatial relationships for a minimum duration. Flock detection thus involves the extraction of spatial features—such as Euclidean inter-pedestrian distance, relative velocity magnitude, differences in motion and facing angles, and trajectory similarity metrics (e.g., Dynamic Time Warping)—and temporal features—such as absolute timestamp differences, duration of spatial cohesion, and stability of motion direction. These spatio-temporal descriptors collectively enable the model to discriminate between coordinated group movement and incidental co-location arising from independent pedestrian motions.
@@ 3.2.1 Time-Bin Classification
$\text{time\_bin}=\left\lfloor\frac{\text{TIMESTAMP}-\text{min}(\text{TIMESTAMP})}{T}\right\rfloor.$ (3)
@@ 3.3.1 Input and Output
3.3.2 Flock Detection Workflow
The flock identification operates as follows:
@@ 4.1.1 Dataset
The group files contain annotations for the groups present on a given day. Only pedestrians in groups are listed, while those walking alone are excluded. Each row represents a single tracked pedestrian within a group and includes the following fields (space-separated values): PEDESTRIANIDPEDESTRIAN-ID, GROUPSIZEGROUP-SIZE, PARTNERID1PARTNER-ID-1… (list of IDs of all other pedestrians in the group), NUMBEROFINTERACTINGPARTNERSNUMBER-OF-INTERACTING-PARTNERS, and INTERACTINGPARTNERID1INTERACTING-PARTNER-ID-1… (list of all socially interacting partners).
For the binary classification model training, we select only data where GROUPSIZEGROUP-SIZE = 2, focusing on pairs of pedestrians from a single day. The pre-trained model is then tested on data from different days, ensuring that the model is evaluated on entirely unseen data.
@@ 4.1.2 Deep Learning Model Preparation
In accordance with the definition of a flock, we extract features that capture spatial and behavioral interactions between pedestrians. These include the inter-distance between pedestrians, and absolute differences in motion properties such as velocity, motion angle, and facing angle. Additionally, we incorporate trajectory-based similarity metrics like Dynamic Time Warping (DTW) to enhance the model’s ability to distinguish between temporary co-movement and sustained flocking behavior.
@@
Feature Scaling: To ensure numerical stability and improve model convergence, each feature is normalized using a specific scaler based on its distribution and sensitivity to outliers:
  • RobustScaler for interDistance: Reduces the influence of extreme values by scaling based on the median and interquartile range.
  • MinMaxScaler for timeDifference: Normalizes values to a fixed range [0, 1], preserving temporal proportionality.
  • StandardScaler for velocityDifference, motionAngleDifference, faceAngleDifference, and dtwValues: Centers features by removing the mean and scaling to unit variance, which is suitable for normally distributed features.
@@ 4.1.3 Multi-agent group data preparation
These bins serve as the basis for further analysis, including the detection of flocking behavior and the execution of pairwise interactions, which are crucial for understanding collective pedestrian dynamics.
@@
(a) Sample Bin Containing Predominantly Groups.
Figure 3: Overview of Sample Bins Containing Group and Single Pedestrians.
In Figure 3, the sample bins illustrate the distribution of group and single pedestrians used as input for the flock detection process. Each bin contains trajectories segmented over a given time window, and the algorithm operates by performing pairwise comparisons within these bins to evaluate spatial and temporal proximity. Based on these comparisons, the model predicts whether each pedestrian pair forms part of a flock, subsequently aggregating positive pairwise predictions to identify complete flocks. This representation highlights both mixed cases—where groups and singles coexist—and bins dominated by either few or many group pedestrians.
@@ 4.3.1 Binary Classification Accuracy by Data Length and Models
Figure 4: Sequential models prediction accuracy for different sequences
@@
4.3.2 Model Training Time Comparison
@@
The runtime comparison results, shown in Fig. 5, indicate that the RNN model consistently achieves the shortest training times across all tested sequence lengths, making it the most computationally efficient among the three architectures. This performance advantage is largely attributable to its relatively simple recurrent structure, which requires fewer parameters and operations per time step.
In contrast, the LSTM model exhibits the longest training times for all sequence lengths, with a peak runtime of 57.1 minutes at a sequence length of 150. This overhead is likely caused by the additional gating and memory cell computations inherent in LSTM’s design, which, while beneficial for long-term dependency modeling, substantially increase computational complexity. The Transformer model’s runtimes are generally shorter than LSTM’s but remain higher than RNN’s, reflecting its reliance on multi-head attention and feed-forward blocks, which scale with sequence length.
An additional observation is that training time does not monotonically increase with sequence length. Instead, the fastest runtimes occur at sequence lengths of 100 and 300 across all models, while the longest runtimes are observed at sequence length 150. This suggests the presence of computational inefficiencies at certain sequence lengths, potentially arising from hardware-level memory allocation or parallelization behavior during training.
@@
  • RNN is the fastest model across all sequence lengths due to its simpler architecture and lower parameter count.
  • LSTM incurs the highest training cost, peaking at 57.1 minutes for sequence length 150, owing to its complex gating mechanisms.
  • Transformer runtimes are generally shorter than LSTM but longer than RNN, reflecting its attention-based design.
  • Sequence lengths of 100 and 300 yield the shortest runtimes across all models, while 150 results in the longest runtimes.
These findings highlight a trade-off between model complexity and training efficiency. While LSTMs and Transformers are generally more expressive for sequence learning tasks, they require significantly longer training times compared to RNNs. Furthermore, sequence length selection can influence training efficiency in non-linear ways, suggesting that computational profiling should be considered alongside predictive performance when choosing model configurations.
@@ 4.3.6 Sample Screens for flock detection
(c) Detected flock size with 6.
Figure 8: Sample Flock Detection Results for Four Different Flock Sizes.
@@ 5 Discussion
In this section, we discuss the results of our flock detection method and highlight key observations regarding the impact of data sequence length and the dependency of features used in the model.
@@
The current feature set for the binary classification model, consisting of inter-agent distance, absolute values of the timestamp, velocity, motion angle, and face angle, has shown promising results in detecting pairs. With a total of 6 features, the model achieved high performance. Interestingly, in some instances, combination of the different features still yielded similar accuracy results. This suggests that the model can effectively detect pairs based on more dynamic features, such as inter-agent distance, velocity, and motion angles. The less significant contribution of static features like center coordinates in certain instances could inform future efforts to optimize the feature set, potentially simplifying the model without sacrificing performance. In practice, reducing the feature space could improve computational efficiency while maintaining high accuracy, especially for real-time applications.